OpenStudy (anonymous):

What are the solutions to the equation x² = 256? Choose all answers that are correct. A. 128 B. –128 C. 16 D. –16

OpenStudy (anonymous):

You have to make root square from both sides

OpenStudy (anonymous):

And you will see what happens next

OpenStudy (anonymous):

\[x^2-256=0\] \[a^2-b^2=\left( a+b \right)\left( a-b \right)\]

OpenStudy (anonymous):

its easy